OpenStudy (anonymous):

how would you integrate (x^2 + 4)/x ?

OpenStudy (anonymous):

\[(x ^{2} + 4)/x = x ^{2}/x + 4/x = x + 4/x \] This should make it more obvious.

OpenStudy (anonymous):

... well then, haha. thanks. thinking WAY too much into it.

OpenStudy (anonymous):

You should get: \[x ^{2}/2 + 4\ln x + C \] as the answer.

OpenStudy (anonymous):

yup- that's what i got. how would i go about solving for x then if C = 1 and the equation = 0? i can't figure out how to find some kind of like term to get a u term alone on the left side